OpenStudy (anonymous):

simplify sqrt -50 - sqrt -16

OpenStudy (anonymous):

= [5sqrt(2) -4] i

OpenStudy (anonymous):

5 i sqrt(2)-4 i

OpenStudy (anonymous):

sqrt -50 + sqrt -16

OpenStudy (lalaly):

\[\sqrt{-50} - \sqrt{-16} \] \[\sqrt{-50} = \sqrt{25} \times \sqrt2 \times \sqrt{-1}\]\[\sqrt{-50}=5\sqrt2 i\] \[\sqrt{-16}= \sqrt{16} \times \sqrt{-1}= 4i\] \[\sqrt{-50} - \sqrt{-16} = 5\sqrt2 i -4i\]

OpenStudy (lalaly):

\[\sqrt{-50}+\sqrt{-16}\]\[=5\sqrt2i + 4i\]